OpenStudy (anonymous):

Can someone help with this?: Find two numbers whose sum is 30, such that the sum of the square of one number plus ten times the other number is a minimum.

OpenStudy (anonymous):

set up a couple equations. x+y=30 \[10x+\sqrt{y}\]=f(x,y) What math is this for?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

first off, I read that wrong.

OpenStudy (anonymous):

sorry about that. Anyway, the equation should really be 10x+y^2->minimum

OpenStudy (anonymous):

so how do you find the two numbers?

OpenStudy (anonymous):

Sorry, I forget how to do it with what you know in algebra, but let's see if I can figure it out as I go along. Using those equations, you also know that 30-y=x and can substitute that in.

OpenStudy (anonymous):

yes I see so far

OpenStudy (anonymous):

so 10(30-y)+y^2 needs to be minimized.

OpenStudy (anonymous):

which means that 300-10y+y^2 needs to be minimized.

OpenStudy (anonymous):

ok so the next step would be setting it to zero?

OpenStudy (anonymous):

I guess. Sorry, I don't remember how to do this the algebra way.

OpenStudy (anonymous):

well I don't really get anywhere doing that anyway. How would you do it?

OpenStudy (anonymous):

do the numbers have to be integers? I would use lagrange multipliers, which I'm pretty sure you haven't learned about, and you don't need for the problem.

OpenStudy (anonymous):

i believe integers

OpenStudy (anonymous):

well to minimize them, 10x would have to be as close to y^2 as possible, so you can just guess and check if they have to be integers to see what gets you the closest.

OpenStudy (anonymous):

10 and 20?

OpenStudy (anonymous):

I don't know. Try asking the question again to see if anyone can do it better.

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

actually, using my way, which I'm not going to explain because it's not the way you're supposed to do it, I got 25 and 5

OpenStudy (anonymous):

where 25 is the one you multiply by 10.

OpenStudy (anonymous):

25 and 5? ok thank you!

OpenStudy (anonymous):

yeah, and you can check that by seeing that no other numbers get you a higher number when you multiply one by 10 and square the other, but I'm almost positive that the answer is 25 and 5.